# Information Transport in Classic-Quantum Hybrid System

### v1

@@ Abstract
Many important quantities in quantum information science, such as entropy and entanglement, are non-linear functions of the density matrix and cannot be expressed as operator observables. Standard open-system approaches evolve only a single copy of the density matrix, making it impossible to track the dynamics of such quantities. A formalism proposed by some of the present authors addressed this challenge by evolving multiple virtual replicas, but was previously limited to the weak-coupling regime. Here, we extend this approach to strong coupling between a quantum system and a classical environment. The resulting multi-replica master equation enables direct evaluation of entropy flow and related metrics in strongly hybridized quantum–classical systems. Our results show that quantum coherence and hybridization jointly suppress net entropy transfer, creating a thermodynamic bottleneck. This framework provides a general tool for studying entropy dynamics and guiding the design of more robust, resource-efficient quantum hardware.
@@ I Introduction
Entropy bridges quantum information science and quantum thermodynamics, serving both as a benchmark for quantum information resources and a limit on thermodynamic efficiency, a key measure for quantum technology such as quantum computers and ultra-efficient heat engines. [1, 2, 3, 4]. In quantum heat engines, entropy governs operating principles and ultimately constrains efficiency [5, 6, 7, 8]. A precise understanding of how entropy is produced is critical to unifying the thermodynamic and information-theoretic descriptions of quantum devices [9, 10, 11, 12, 13, 14, 15, 16]. Real-time entropy measurement cannot take place by physical observables–such as energy, charge, and spin–because the nonlinearity of the density matrix in the definition of entropy makes it a non-observable quantity. However, entropy production can be tracked in the statistics of physical currents [17, 18].
@@
Subsequent developments established how to measure entropy by R/FCS correspondence–an exact correspondence between entropy flow and the full statistics of energy transfer  [30, 31, 18]. This discovery demonstrated how quantum superposition can impact the statistics of energy transfer corresponding to the coherent entropy flow. The emergence of a superposition-dependent entropy flow and its direct connection to the complex quantum statistics of energy transfer has the potential to heavily affect a significant advancement for quantum circuit theory [32, 33, 34], by providing crucial insights for accurately modeling quantum nanostructures [35, 36, 37] and controlling the flow of information and physical quantities in them [38, 39, 33, 40].
@@ II Weak Coupling Formalism
In a closed system the Rényi entropy is conserved. Observing only a part of a composite system converts this global conservation law into a current of Rényi entropy that flows between subsystems [23]. We study a quantum system that mediates information between several reservoirs. A reservoir acting as a classical probe, labeled by the index $b$, is a classical environment held at a fixed, low temperature and is coupled to the system through a small transition rate $\Gamma_{b}$ that weakly perturbs the quantum system. For the purpose of evaluating the entropy exchanged with the probe, we work in the weak–coupling regime $\Gamma_{b}\ll\Gamma_{e}$; here $\Gamma_{e}$ collectively denotes the couplings to all other reservoirs $e$. All reservoirs are maintained in the linear–response regime: they remain at fixed temperature, do not interact directly with each other, and any nonequilibrium effects arise solely within the quantum node as a shared small subsystem between them. The reduced dynamics under the couplings $\Gamma_{e}$ are obtained by solving an appropriate open quantum system approach and tracing out environments. The resulting density matrix governs the time evolution of the quantum node $\rho(t)$, which is then substituted into Eq. (3) to compute the Rényi entropy current flowing into the probe reservoir $b$.
Since the total Rényi entropy of the composite system is conserved, the resulting entropy current depends only on the instantaneous state of the central subsystem, which therefore acts as an entropy ‘valve.’ Evaluating the flow of entropy from Eq. (3) requires a replicated quantum master equation. More precisely, $M$ replicas of the closed system are considered: for both the quantum system and the probe reservoir, separate closed-timeline Keldysh-type contours are considered in each replica with backward (forward) flow for bra (ket) states. Since the measurement takes place in the probe reservoir, the contour associated with the quantum system gets closed in each replica to denote the tracing of the quantum system out, leaving only the probe reservoir branches connected across replicas. This results in the quantum master equation for a multi-replicated density matrix of probe reservoir, which after proper final tracing, will result in the entropy flow to probe reservoir.
@@
In the absence of stationary quantum coherence, the expressions are reduced to standard energy-exchange statistics. Persistent coherence can make the net flow negative, a clear signature of genuinely quantum back-action among the replicas.
Example 1: Consider a simple quantum system consisting of a qubit with two states, $|0\rangle$ and $|1\rangle$, separated by energy $\hbar\omega=(E_{1}-E_{0})$. This qubit is driven by an oscillating external field described by the Hamiltonian $H_{dr}=\hbar\sum_{m,n=\{0,1\}}\Omega|m\rangle\langle n|\exp(-i\omega_{dr}t)$, with driving frequency $\omega_{\text{dr}}$ close to the qubit’s resonance frequency $\omega$ and amplitude $\hbar\Omega$. The qubit mediates energy and information exchange between two reservoirs: a weakly coupled probe reservoir at zero temperature and a stronger-coupled environment at finite temperature $T_{b}$.
@@
Ref. [24] shows that the stationary-state Rényi entropy flow into the probe reservoir comprises two distinct parts: the incoherent (probabilistic) contribution $F_{M}^{\text{incoh}}=G_{M}\omega(\Gamma^{b}_{\downarrow}p_{1}-\Gamma^{b}_{\uparrow}p_{0})$, and the coherent contribution $F_{M}^{\text{coh}}=G_{M}\omega\Gamma^{b}|\rho_{01}|^{2}$, with $p_{0,1}$ representing the populations and $\rho_{01}$ the coherence of the qubit density matrix in the energy basis; $\Gamma^{b}\equiv\Gamma^{b}_{\downarrow}-\Gamma^{b}_{\uparrow}$. The prefactor $G_{M}$ is defined as $G_{M}\equiv Mn_{\text{B}}(\Omega_{M}/T)/[n_{\text{B}}(\Omega_{M-1}/T)n_{\text{B}}(\Omega_{1}/T)]$. In the limit of $M\to 1$ one can recovers the von Neumann entropy for this system which turns out to still have the coherent flow persistent in it:
$F_{vN}=\left(\Gamma^{b}_{\downarrow}p_{1}-\Gamma^{b}_{\uparrow}p_{0}-\Gamma^{b}|\rho_{01}|^{2}\right)\omega/T_{\rm{b}}$ (5)
In the presence of stationary quantum coherence in the system, the total entropy flow is reduced but remains non-negative, indicating that the probe functions as an entropy sink, with qubit coherence merely slowing the exchange rate.
@@
$F_{vN}=[n_{\text{B}}(\omega/T_{\text{h}})-n_{\text{B}}(\omega/T_{b})]\chi^{b}\omega/T_{\text{b}}$
Notably, the entropy flow depends solely on the oscillator and reservoir temperatures and remains unaffected by the external drive amplitude. The flow reverses sign at $T_{b}=T_{\text{h}}$, becoming negative when the quantum oscillator coherently injects entropy into the colder reservoir.
The examples discussed here and others such as photovoltaic cells [30, 31] illustrate that a consistent extension of information measures to quantum heat engines is feasible and can serve for pathfinding the next generation of quantum heat engines.
@@ III Quantum-Classical Hybridization
In a composite system consisting of a number of classical systems coupled to a mediating quantum system, the textbook Hamiltonian is split as $H=H_{Q}+\sum_{C}H_{C}+\sum_{C}H_{\mathrm{QC}}$, with Quantum-Classic interaction $\lVert H_{\mathrm{QC}}\rVert$ much smaller than the minimum energy gaps in the quantum Hamiltonian $H_{Q}$ and the classical systems $H_{C}$. One diagonalizes $H_{Q}$ and $H_{C}$ separately and treats $H_{\mathrm{QC}}$ perturbatively. Each reservoir, assumed macroscopic with a continuous spectrum, sits at its fixed temperature, introducing only small corrections to the isolated dynamics.
@@
—We consider a device in which the quantum subsystem is strongly correlated with its classical probing environment. Here we consider the example of a low temperature classical system that serves as a probe measurement with transition rate $\Gamma_{b}$, and a second classical environment at finite temperature with transition rate $\Gamma_{e}$. Throughout, sometimes we use the unindex $\Gamma$ instead of $\Gamma_{e}$.
Strong coupling merges classical and quantum dynamics, so that the usual system-bath separation no longer applies. Accurate descriptions must therefore include the hybridized degrees of freedom explicitly and compute transport and feedback. In Fig. (1) the classical systems (2) and (3) are strongly coupled so that treating them as distinct subsystems is inaccurate. In other words, estimating the subsystem energy spectrum in the absence and in the presence of other strongly-interacting subsystems will be completely different so that one will not be an approximation of the other one. What we present below is going to emphasize this distinction by comparing weak and strong coupling strengths.
@@
—The Rényi entropy in Eq. (2) implicitly assumes that the $M$ replica density matrices are separable, sharing only the minimal degrees of freedom. Whether among atomic orbitals in chemistry [51], phonon–photon polaritons in solids [52], or spin–orbit coupling in atomic systems [53],, hybridization is the coherent mixing of basis states that were once subsystem-specific. Therefore, we work in a tensor product basis that spans the entire Hilbert space. The Hamiltonian takes a block structure (often sparse): diagonal blocks hold the isolated energies, while off-diagonal blocks encode the couplings that drive hybridization. If these couplings dominate, the ensuing entanglement is so widespread that states can no longer be unambiguously assigned to any single component.
When three quantum subsystems C, Q, and $\mathrm{\overline{C}}$ are hybridized, their joint state resides in the composite Hilbert space and is fully characterized by a density operator $\rho\in\mathrm{L}(\mathcal{H}_{C}\otimes\mathcal{H}_{Q}\otimes\mathcal{H}_{\overline{C}})$, with $\mathrm{L}$ being the linear map within the Hilbert space.
@@
—Combining the concept of tensor-network time evolution with the multi-replica framework enables a redefinition of the Rényi entropy for hybridized systems. In this context, the notion of replica separability from the original definition of Rényi entropy can be generalized, while still yielding a scalar quantity. As a reminder, consider a probe reservoir $\overline{\mathrm{C}}$ and a quantum system $Q$ coupled to an environment $C$. Within the KAN formalism, the probe entropy is obtained from
$\mathrm{Tr}_{\overline{C}}\left(\rho_{\overline{C}}\right)^{M},$ (7)
which implies that all contours associated with $Q$ and $C$ are traced out within each replica, except for the probe density matrix. The latter spans all replicas and is traced only at the final stage.
Between subsystems weak coupling can set a boundary; however, hybridization of subsystems makes the boundary dynamically ill-defined. To address this breakdown, consider the density matrix subject to time evolution $d\rho_{{}_{CQ{\overline{C}}}}/dt=\mathcal{L}\rho_{{}_{CQ{\overline{C}}}}$ with $\mathcal{L}$ being the so-called Liouvillian operator, a typically non-Hermitian operator to be determined perturbatively or using renormalization group methods.
@@
Let us now define the generalized replicated density matrix for $M$ hybridized quantum systems only, after tracing out environment in each replia, the probe reservoir and the quantum system go together in all $M$ replicas. This allows not only exchanges of particles in one replica to take place, but also exchanges in between replicas, similar to the weak coupling regime. By tracing out the probe reservoir the we are left with $M$ replicas of the quantum system density matrix, which we denote as
@@
which defines Rényi entropy in strong coupling regime for convoluted systems so that they cannot be identified as isolated systems anymore.
@@ III.1 Evolution of Generalized Density Matrix
Computing the Rényi entropy flow $dS_{M}/dt$ in strongly coupled systems from Eq. (10) requires a novel extension of the KAN diagrams in order to represent $d\hat{R}_{M}/dt$, i.e. the time evolution of the generalized density matrix defined in Eq. (9).
@@ III.2 Diagrammatic Keldysh Contours
Motivated by the definition of the Rényi entropy through the generalised density matrices in Eqs. (9) and (10), we can extend the KAN formalism from the weak- to the strong-coupling regime. For this, we note that the final result in Eq. (14) shows that the Rényi entropy flow does not require tracing out all contours; instead it is fixed by the lowest non-zero eigenvalue of the Liouvillian operator.
For a classic–quantum–classic hybridised system, illustrated in Fig. 2, we consider three contours in each replica. Time runs from the distant past at the left to the present on the right. Every contour first moves backwards in time, representing a bra state $\langle\cdot|$, passes through the initial density matrix (depicted as a box with label $\rho_{-\infty}$), and then returns forwards in time toward the present to represent the corresponding ket state $|\cdot\rangle$ of that subsystem.
@@ IV A Hybridized Qubit As Energy Valve
We will now apply the strong coupling formalism described above to a simple quantum system, which consists of an externally driven qubit. First let us describe the quantum system part. Let us consider the qubit frequency is $\hbar\omega=E_{1}-E_{0}$ and externally driven by amplitude $\hbar\Omega$ and the frequency $\omega_{\mathrm{dr}}$, making the Hamiltonian $\hat{H}_{dr}(t)=\hbar\Omega\cos(\omega_{dr}t)\left(|0\rangle\langle 1|+H.c.\right)$. The driving amplitude $\Omega$ is not strong enough to inject new photon into the system, i.e. $\Omega\ll\omega$.
In order to make the driving part independent explicitly time-dependent, we transform the Hamiltonian into a frame rotating with the driving frequency. Performing the rotating wave approximation (RWA) the Hamiltonian becomes
@@
Two reservoirs $e$ (environment) and $b$ (probe) are coupled to the quantum system. Linear response theory is considered to let the reservoirs interact only with the qubit. Each reservoir $R=\{e,b\}$ carries the free Hamiltonian $\hat{H}_{R}=\sum_{k}\omega_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}$, with $\hat{a}_{k}$ denoting the ladder annihilation operator in the reservoir $R$ and the frequency mode $k$. In other words, we consider each reservoir to be represented by a complete set of oscillators at a continuous frequency. The spectrum of reservoir frequencies $\omega_{k}$ is assumed to be sufficiently dense to be considered continuous.
The quantum system is coupled to the reservoirs R by individual oscillators in it by the following interaction Hamiltonian $H_{QR}=\sum_{ij}|i\rangle\langle j|\hat{X}^{(R)}_{ij}$, with $\hat{X}_{ij}^{(R)}$ being the operators in the space of reservoir and defined as $\hat{X}^{(R)}_{ij}=\sum_{k}c^{(R)}_{ij,k}\hat{a}_{k}e^{-i\omega_{k}t}+H.c.$ The operator $\hat{a}_{k}$ being annihilation operator of mode frequency $\omega_{k}$. The coupling strength between the $0\leftrightarrow 1$ transition in the qubit and $i\leftrightarrow j$ transition of the mode $k$ determines the complex number $c^{(R)}_{ij,k}$. In fact this coefficient is an important quantity as it determines the strength of the coupling between qubit and reservoir as shown by connectors of their degrees of freedom in Fig. (1). Throughout, we consider driving frequency to match with the qubit frequency, i.e. $\delta=0$. Rotating the interaction of quantum system with each classical system of reservoir transforms to the following
@@
—The dynamical (retarded) susceptibility $\chi_{AB}(\omega)$ quantifies how fast and strong an observable $A$ under external driving field $f(t)$ is perturbed into $\delta\langle A(t)\rangle$ as the result of the influence of another observable $B$ at earlier time. In the linear-response theory, this can be written as $\delta\langle A(t)\rangle=\!\int_{-\infty}^{\infty}\!dt^{\prime}\,\chi_{AB}(t-t^{\prime})\,f(t^{\prime})$ with the dynamical susceptibility being $\chi_{AB}(t)=\frac{i}{\hbar}\,\Theta(t)\,\bigl{\langle}[A(t),B(0)]\bigr{\rangle}_{\!\mathrm{eq}}$ with the index $eq$ indication the equilibrium fluctuations.
The dynamical susceptibility is related to the correlators of $\hat{X}^{(R)}$ (for brevity here we drop the index $(R)$) defined as $S_{ij,pq}(\tau)\equiv\textup{Tr}_{R}\{\hat{X}_{ij}(0)\hat{X}_{pq}(\tau)\rho_{R}\}_{\!\mathrm{eq}}$. Note that the correlator depends on the time delay $\tau$ between the two operators $\hat{X}$ acting in the reservoir, with the time delay being minimally zero. We need to evaluate the half-sided Fourier transform of the correlator between the two intervals: $\int_{0}^{\infty}d\tau\exp({i\omega\tau})S_{ij,pq}(\pm\tau)\equiv\frac{1}{2}S_{ij,pq}(\pm\omega)\pm i\Pi_{ij,pq}(\pm\omega)$, which plays a crucial role in the evaluation, especially due to its application in the Kubo-Martin-Schwinger [50] and its generalization [24] that relates the correlator to the dynamical susceptibilities $\tilde{\chi}_{ij,pq}(\nu)$ of the reservoir:

### v2

# Information Transport in Classical-Quantum Hybrid System

@@ Abstract
Many of the quantities that matter in quantum technology, e.g. entropy, entanglement, etc., cannot be associated with operator observables, because they depend non-linearly on the quantum state’s density matrix. The non-linearity provokes another bigger issue: Standard open-system equations evolve just one copy of that matrix, so we cannot track how such elusive quantities change. A recent formalism by Ansari and Nazarov sidestepped this by evolving multiple virtual replicas at once, but only in the weak-coupling regime. Here, we remove that limitation. Our generalized, multi-replicative master equation follows entropy flow and other vital metrics even when a quantum device interacts strongly with a classical environment. We show quantum coherence and hybridization jointly act to inhibit the net transfer of entropy, effectively introducing a “bottleneck” in the thermodynamic process. This sharper view of quantum–classical hybridization points the way toward more robust, resource-aware quantum hardware.
@@ I Introduction
Entropy bridges quantum information science and quantum thermodynamics, serving both as a benchmark for quantum information resources and a limit on thermodynamic efficiency, a key measure for quantum technology such as quantum computers and ultra-efficient heat engines. [1, 2, 3, 4]. In quantum heat engines, entropy governs operating principles and ultimately constrains efficiency [5, 6, 7, 8]. A precise understanding of how entropy is produced is critical to unifying the thermodynamic and information-theoretic descriptions of quantum devices [9, 10, 11, 12, 13, 14, 15, 16]. Real-time entropy measurement cannot take place by physical observables–such as energy, charge, and spin–because of the nonlinearity of the density matrix in the definition of entropy, makes it a non-observable quantity. However, entropy production can be tracked in the statistics of physical currents [17, 18].
@@
Subsequent developments established how to measure entropy by R/FCS correspondence–an exact correspondence between entropy flow and the full statistics of energy transfer  [30, 31, 18]. This discovery demonstrated how quantum superposition can impact the statistics of energy transfer corresponding to the coherent entropy flow. The emergence of a superposition-dependent entropy flow and its direct connection to the complex quantum statistics of energy transfer has the potential to heavily affect a significant advancement for quantum circuit theory [32, 33, 34, 35, 36], by providing crucial insights for accurately modeling quantum nanostructures [37, 38, 39, 40, 41] and controlling the flow of information and physical quantities in them [42, 43, 34, 44, 45, 46].
@@ II Weak Coupling Formalism
In a closed system the Rényi entropy is conserved. Observing only a part of a composite system converts this global conservation law into a current of Rényi entropy that flows between subsystems [23]. We study a quantum system that mediates information between several reservoirs. A reservoir acting as a classical probe, labeled by the index $b$, is a classical environment held at a fixed, low temperature and is coupled to the system through a small transition rate $\Gamma_{b}$ that weakly perturbs the quantum system [56]. For the purpose of evaluating the entropy exchanged with the probe, we work in the weak–coupling regime $\Gamma_{b}\ll\Gamma_{e}$; here $\Gamma_{e}$ collectively denotes the couplings to all other reservoirs $e$. All reservoirs are maintained in the linear–response regime: they remain at fixed temperature, do not interact directly with each other, and any nonequilibrium effects arise solely within the quantum node as a shared small subsystem between them. The reduced dynamics under the couplings $\Gamma_{e}$ are obtained by solving an appropriate open quantum system approach and tracing out every environment. The resulting density matrix governs the time evolution of the quantum node $\rho(t)$, which is then substituted into Eq. (3) to compute the R ’enyi entropy current flowing into the probe reservoir $b$.
Since the total Rényi entropy of the composite system is conserved, the resulting entropy current depends only on the instantaneous state of the central subsystem, which therefore acts as an entropy “valve.” Evaluating the flow of entropy from Eq. (3) requires a replicated quantum master equation. More precisely, $M$ replicas of the closed system are considered: for both the quantum system and the probe reservoir, separate closed-timeline Keldysh-type contours are considered in each replica with backward (forward) flow for bra (ket) states. Since the measurement takes place in the probe reservoir, the contour associated with the quantum system gets closed in each replica to denote the tracing of the quantum system out, leaving only the probe reservoir branches connected across replicas. This results in the quantum master equation for a multi-replicated density matrix of probe reservoir, which after proper final tracing, will result in the entropy flow to probe reservoir.
@@
In the absence of stationary quantum coherence, the expression are reduced to standard energy-exchange statistics. Persistent coherence can make the net flow negative, a clear signature of genuinely quantum back-action among the replicas.
Example 1: Consider a simple quantum system consisting of a qubit with two states, $|0\rangle$ and $|1\rangle$, separated by energy $\hbar\omega=(E_{1}-E_{0})$. This qubit is driven by an oscillating external field described by the Hamiltonian $H_{dr}=\hbar\sum_{m,n=\{0,1\}}\Omega|m\rangle\langle n|\exp(-i\omega_{dr}t)+H.c.$, with $m\neq n$ and the driving frequency $\omega_{\text{dr}}$ close to the qubit’s resonance frequency $\omega$ and amplitude $\hbar\Omega$. The qubit mediates energy and information exchange between two reservoirs: a weakly coupled probe reservoir at nearly zero temperature and a stronger-coupled environment at finite temperature $T_{b}$.
@@
Ref. [24] shows that the stationary-state Rényi entropy flow into the probe reservoir comprises two distinct parts: the incoherent (probabilistic) contribution $F_{M}^{\text{incoh}}=G_{M}\omega(\Gamma^{b}_{\downarrow}p_{1}-\Gamma^{b}_{\uparrow}p_{0})$, and the coherent contribution $F_{M}^{\text{coh}}=G_{M}\omega\Gamma^{b}|\rho_{01}|^{2}$, with $p_{0,1}$ representing the populations and $\rho_{01}$ the coherence of the qubit density matrix in the energy basis; $\Gamma^{b}\equiv\Gamma^{b}_{\downarrow}-\Gamma^{b}_{\uparrow}$. The prefactor $G_{M}$ is defined as $G_{M}\equiv Mn_{\text{B}}(\Omega_{M}/T)/[n_{\text{B}}(\Omega_{M-1}/T)n_{\text{B}}(\Omega_{1}/T)]$. In the limit of $M\to 1$ one recovers the von Neumann entropy for this system which turns out to still have the coherent flow persistent in it:
$F_{\rm{vN}}=(\Gamma^{b}_{\downarrow}p_{1}-\Gamma^{b}_{\uparrow}p_{0}-\Gamma^{b}|\rho_{01}|^{2})\omega/T_{\rm{b}}$ (5)
In the presence of stationary quantum coherence in the system, the toral entropy flow slows down but cannot become negative, which indicates that the probe is entropy sink, yet the qubit coherence slows down the exchange rate.
@@
$F_{\text{vN}}=[n_{\text{B}}(\omega/T_{\text{h}})-n_{\text{B}}(\omega/T_{b})]\chi^{b}\omega/T_{\text{b}}$
Notably, the entropy flow depends solely on the oscillator and reservoir temperatures and remains unaffected by the external drive amplitude. The flow reverses sign at $T_{b}\geq T_{\text{h}}$, becoming negative when the quantum oscillator coherently injects entropy into the colder reservoir.
The examples discussed here and others such as photovoltaic cells [30, 31] illustrate that a consistent extension of information measures to quantum heat engines is feasible and can serve for pathfinding next generation of quantum heat engines.
@@ III Quantum-Classical Hybridization
In a composite system consisting of a number of classical system coupled to a mediating quantum system, the textbook Hamiltonian is split as $H=H_{Q}+\sum_{C}H_{C}+\sum_{C}H_{\mathrm{QC}}$, with Quantum-Classical interaction $\lVert H_{\mathrm{QC}}\rVert$ much smaller than the minimum energy gaps in the quantum Hamiltonian $H_{Q}$ and the classical systems $H_{C}$. One diagonalizes $H_{Q}$ and $H_{C}$ separately and treats $H_{\mathrm{QC}}$ perturbatively. Each reservoir, assumed macroscopic with a continuous spectrum, sits at its fixed temperature, introducing only small corrections to the isolated dynamics.
@@
—We consider a device in which the quantum subsystem is strongly correlated with its classical probing environment. Here we consider the example of a low temperature classical system that serves as a probe measurement with transition rate $\Gamma_{b}$, and a second classical environment at finite temperature with transition rate $\Gamma_{e}$.
Strong coupling merges classical and quantum dynamics, so that the usual system-bath separation no longer applies. Accurate descriptions must therefore include the hybridized degrees of freedom explicitly and compute transport and feedback. In Fig. (1) the classical systems (2) and (3) are strongly coupled so that treating them as distinct subsystems is inaccurate. In other words, estimating the subsystem energy spectrum in the absence and in the presence of other strongly-interacting subsystems will be completely different so that one will not be an approximation of the other one. What we present below is going to emphasize on the this distinction by comparing weak and strong coupling strengths.
@@
—The Rényi entropy in Eq. (2) implicitly assumes that the $M$ replica density matrices are separable, sharing only the minimal degrees of freedom. Whether among atomic orbitals in chemistry [58], phonon–photon polaritons in solids [59], or spin–orbit coupling in atomic systems [60],, hybridization is the coherent mixing of basis states that were once subsystem-specific. Therefore, we work in a tensor product basis that spans entire Hilbert space. The Hamiltonian takes a block-structured (often sparse): diagonal blocks hold the isolated energies, while off-diagonal blocks encode the couplings that drive hybridization. If these couplings dominate, the ensuing entanglement is so widespread that states can no longer be unambiguously assigned to any single component.
When three quantum subsystems C, Q, and $\mathrm{\overline{C}}$ are hybridized, their joint state resides in the composite Hilbert space and is fully characterized by a density operator $\rho\in\mathrm{L}(\mathcal{H}_{C}\otimes\mathcal{H}_{Q}\otimes\mathcal{H}_{\overline{C}})$, with $\mathrm{L}$ being the linear map within the Hilbert space to the density matrix representation.
@@
[…]
Between subsystems weak coupling can set a boundary; however, hybridization subsystems makes the boundary dynamically ill-defined. To address this breakdown, consider the density matrix subject to time evolution of the quantum system and a partial subset of coupled classical system, e.g. $d\rho_{{}_{Q{\overline{C}}}}/dt=\mathcal{L}\rho_{{}_{Q{\overline{C}}}}$ with $\mathcal{L}$ being the so-called Liuovillian operator, a typically non-Hermitian operator to be determined perturbatively or using renormalization group methods.
@@
Let us now define the generalized replicated density matrix for $M$ hybridized quantum systems only, after tracing out environment in each replia, the probe reservoir and the quantum system go together in all $M$ replicas. This allows not only exchanges of particles in one replica to take place, but also exchanges in between replicas becomes a possibility, similar to weak coupling regime. By tracing out the probe reservoir we are left with $M$ replicas of the quantum system density matrix, which we denote as
@@
which defines Rényi entropy in strong coupling regime for convoluted systems so that they cannot be identified as isolated system anymore.
@@ III.1 Evolution of Generalized Density Matrix
Computing the Rényi entropy flow $dS_{M}/dt$ in strongly coupled systems from Eq. (9) requires a novel extension of the KAN diagrams in order to represent $d\hat{R}_{M}/dt$, i.e. the time evolution of the generalized density matrix defined Eq. (8).
@@ III.2 Diagrammatic Keldysh Contours
Motivated by the definition of the Rényi entropy through the generalised density matrices in Eqs. (8) and (9), we can extend the KAN formalism from the weak- to the strong-coupling regime. For this, we note that the final result in Eq. (13) shows that the Rényi-entropy flow does not require tracing out all contours; instead it is fixed by the lowest non-zero eigenvalue of the Liouvillian operator.
For a classical–quantum–classical hybridised system, illustrated in Fig. 2, we consider three contours in each replica. Time runs from the distant past at the left to the present on the right. Every contour first moves backwards in time, representing a bra state $\langle\cdot|$, passes through the initial density matrix (depicted as a box with label $\rho_{-\infty}$), and then returns forwards in time toward the present to represent the corresponding ket state $|\cdot\rangle$ of that subsystem.
@@ IV A Hybridized Qubit As Energy Valve
We will now apply the strong coupling formalism described above to a simple quantum system, which consists of an externally driven qubit. First let us describe the quantum system part. Let us consider the qubit frequency is $\hbar\omega=E_{1}-E_{0}$ and externally driven by amplitude $\hbar\Omega$ and the frequency $\omega_{\mathrm{dr}}$, making the Hamiltonian $\hat{H}_{dr}(t)=\hbar\Omega\cos(\omega_{dr}t)\left(|0\rangle\langle 1|+H.c.\right)$. The driving amplitude $\Omega$ is not that strong to inject new photon into the system, i.e. $\Omega\ll\omega$.
In order to make the driving part independent explicitly from time, we transform the Hamiltonian into a frame rotating with the driving frequency. Performing the rotating wave approximation (RWA) so that the Hamiltonian of the qubit subsystem becomes
@@
Two reservoirs $e$ (environment) and $b$ (probe) are coupled to the quantum system. Linear response theory is considered to let the reservoirs interact only with the qubit. Each reservoir $R=\{e,b\}$ carries the free Hamiltonian $\hat{H}_{R}=\sum_{k}\omega_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}$, with $\hat{a}_{k}$ denoting the ladder annihilation operator in the reservoir $R$ and the frequency mode $k$. In other words we consider each reservoir can be represented by a complete set of oscillators at a continuous frequency. The spectrum of reservoir frequencies $\omega_{k}$ is assumed to be sufficiently dense to be considered continuous.
The quantum system is coupled to the reservoirs R by individual oscillators in it by the following interaction Hamiltonian $H_{QR}=\sum_{ij}|i\rangle\langle j|\hat{X}^{(R)}_{ij}$, with $\hat{X}_{ij}^{(R)}$ being the operators in the space of reservoir and defined as $\hat{X}^{(R)}_{ij}=\sum_{k}c^{(R)}_{ij,k}\hat{a}_{k}e^{-i\omega_{k}t}+H.c.$ The operator $\hat{a}_{k}$ being annihilation operator of mode frequency $\omega_{k}$. The coupling strength between the $0\leftrightarrow 1$ transition in the qubit and $i\leftrightarrow j$ transition of the mode $k$ determines the complex number $c^{(R)}_{ij,k}$. In fact this coefficient is important quantity as it determines the strength of the coupling between qubit and reservoir as shown by connectors of their degrees of freedom in Fig. (1). Throughout we consider driving frequency matches with the qubit frequency, i.e. $\delta=0$. Rotating the interaction of quantum system with each classicla system of reservoir transforms to the following
@@
—The dynamical (retarded) susceptibility $\chi_{AB}(\omega)$ quantifies how fast and strong an observable $A$ under external driving field $f(t)$ is perturbed into $\delta\langle A(t)\rangle$ as the result of the influence of another observable $B$ at earlier time. In the linear-response theory, this can be written as $\delta\langle A(t)\rangle=\!\int_{-\infty}^{\infty}\!dt^{\prime}\,\chi_{AB}(t-t^{\prime})\,f(t^{\prime})$ with the dynamical susceptibility being $\chi_{AB}(t)=\frac{i}{\hbar}\,\Theta(t)\,\bigl{\langle}[A(t),B(0)]\bigr{\rangle}_{\!\mathrm{eq}}$ with the index $eq$ indicating the equilibrium fluctuations.
The dynamical susceptibility is related to the correlators of $\hat{X}^{(R)}$ (for brevity here we drop the index $(R)$) defined as $S_{ij,pq}(\tau)\equiv\textup{Tr}_{R}\{\hat{X}_{ij}(0)\hat{X}_{pq}(\tau)\rho_{R}\}_{\!\mathrm{eq}}$. Note that the correlator depends on the time delay $\tau$ between the two operators $X$’ acting in the reservoir, with the time delay being minimally zero. We need to evaluate the half-sided Fourier transform of the correlator between the two intervals: $\int_{0}^{\infty}d\tau\exp({i\omega\tau})S_{ij,pq}(\pm\tau)\equiv\frac{1}{2}S_{ij,pq}(\pm\omega)\pm i\Pi_{ij,pq}(\pm\omega)$, which plays a crucial role in the evaluation, especially due to its application in the Kubo-Martin-Schwinger [57] and its generalization [24] that relates the correlator to the dynamical susceptibilities $\tilde{\chi}_{ij,pq}(\nu)$ of the reservoir: